Three members of your family (two parents and yourself) went to buy three cell phones. The phones come in three colors: red, blue and black. You all pick them randomly. What is the probability that all three phones will be red or blue?
Answer by stanbon(75887)

Three members of your family (two parents and yourself) went to buy three cell phones. The phones come in three colors: red, blue and black. You all pick them randomly. What is the probability that all three phones will be red or blue?
----
For each individual: P(red or blue) = P(red) + P(blue) = 1/3 + 1/3 = 2/3
---
For all three indicuals: P(red or blue) = (2/3)^3 = 8/27
===========================
Cheers,
Stan H.